Using 10 mL of 0.2 M H2SO4 solution, calculate the volume of water necessary to create a
0.05 M solution of H2SO4. What is the final volume of the solution?
1
Expert's answer
2020-05-10T14:57:51-0400

CM = n / V

n = CM x V = 0.2 mol/L x 0.01 L = 0.002 mol

V = n / CM = 0.002 / 0.05 = 0.04 L

∆V = 40 - 10 = 30 ml of water